OpenStudy (anonymous):

17 Which of the following properties of a ball is conserved as it falls freely in a vacuum? kinetic energy gravitational potential energy monentum mechanical energy 18 How much work is done when a bucket of mass 1.5 kg with 10 kg of water in it is pulled up from the bottom of a well 8 m deep? Take g=98ms2 100.3 J 2000.5 J 901.6 J 453.7 J 19 A 65 kg sprinter completes a 100 m race in 9.83 s. Calculate the average kinetic energy of the sprinter 3.36 kJ 2.11 kJ 34.15 kJ 65.42 kJ

OpenStudy (ash2326):

I'll answer the first one, Ball is falling freely in vacuum, it has only one force acting,which is gravitational force. So it has a gravitational potential energy, It's in motion so it has a Kinetic energy. As it falls, its gravitational potential energy decreases and kinetic energy increases. But if we sum the two energies, which is known as the Mechanical Energy, it'll be constant. Momentum is always changing,as the velocity is changing. Do you understand this @OLABODEBG?

OpenStudy (anonymous):

Gravitational Potential Energy W =(m1+m2)gh=901.6 K.E=1/2mv^2=1/2*65*100/9.83=3363.383

OpenStudy (anonymous):

1. Mechanical energy is conserved. 2. What you need to calculate is the increase in potential energy when it is brought up by a height of 8m PE = mgh = (1.5 + 10) * 9.8 * 8 = work done 3. avg. speed = distance/time = 100/9.83 = v(say) avg. K.E. = (1/2) * m * v*v